OpenStudy (anonymous):

a^2-2ab-15b^2 factor completly please ^ this means square

OpenStudy (anonymous):

(a+3b)(a-5b)

OpenStudy (anonymous):

\[a^2-5ab+3ab-15b^2\]

OpenStudy (cwtan):

(a-5b)(a+3b) 1*15=15 5*3=15 the most probably set is 3*5 since -5+3=-2 and -5*3=-15 so the solution is (a-5b)(a+3b)